. Determine the effective number of lattice points for a face-centred cubic lattice


1
Expert's answer
2021-01-22T09:56:53-0500

Solution:

"\dfrac{1}{8}" at each of the 8 corners plus "\dfrac{1}{2}" on each of the 6 faces = 1 + 3 = 4.